OpenStudy (anonymous):

The ratio of two numbers is 5 to 3. The sum of the numbers is 16. What are the numbers? 8 and 8 9 and 7 5 and 3 10 and 6

OpenStudy (anonymous):

10 and 6

OpenStudy (anonymous):

a/b = 5/3 - a= (5/3)b a+b = 16 - (5/3)b + b = 16 (8/3)b = 16 b=(16*3)/8=6 so a= 10 what i've done was solving the equation in terms of b :) hope you got the concept